OpenStudy (anonymous):

The y-coordinate of the solution to the system shown is? 5x + y = 6 5x + 3y = -4

OpenStudy (anonymous):

subtract from both equation then we find 5x+y-5x-3y=6-(-4) =>y-3y=6+4 =>-2y=10 =>y=-5

OpenStudy (anonymous):

DUHH GOLD

OpenStudy (anonymous):

?

OpenStudy (anonymous):

gold is my friend and he was looking on here so thats why i said DUHHH GOLD